OpenStudy (anonymous):

Solve for the variable. Show your work. * Remember that a variable cannot be negative. 75 – n = 26

OpenStudy (anonymous):

@sleepyhead314 @Michele_Laino @AlexandervonHumboldt2 Can someone please help me

OpenStudy (jb1625):

This easy

OpenStudy (jb1625):

Do 75-26=?

OpenStudy (jb1625):

Then you have your answer

OpenStudy (jb1625):

Are you there?

OpenStudy (jb1625):

@INeedHelpXoXo123

OpenStudy (jb1625):

Send me a message when you come back

OpenStudy (texaschic101):

75 - n = 26 --- add n to both sides 75 + n - n = 26 + n --- the n's cancel out on the left side 75 = 26 + n -- now subtract 26 from both sides 75 - 26 = n 49 = n

OpenStudy (anonymous):

back sorry i had to make lunch for my brother :/

OpenStudy (jb1625):

lol